A new-born child can be given 1 or 2 names. In how many ways can a child be named if we can choose from 100 names?
Answer by stanbon(75887)

A new-born child can be given 1 or 2 names. In how many ways can a child be named if we can choose from 100 names?
-----------------
100C1 + 100C2 = 100 + (100*99)/(1*2) = 100 + 50*99
=========
Cheers,
Stan H.